# CLARITY: Clinical Assistant for Routing, Inference, and Triage

### v1

@@ Abstract
We present CLARITY (Clinical Assistant for Routing, Inference, and Triage), an AI-driven platform designed to facilitate patient-to-specialist routing, clinical consultations, and severity assessment of patients’ conditions. Its hybrid architecture combines a Finite State Machine (FSM) for structured dialogue flows with collaborative agents that employ Large Language Model (LLM) to analyze symptoms and prioritize referrals to appropriate specialists. Built on a modular microservices framework, CLARITY ensures safe, efficient, and robust performance, flexible and readily scalable to meet the demands of existing workflows and IT solutions in healthcare.
We report integration of our clinical assistant into a large-scale nation-wide inter-hospital IT platform, with over 55,000 content-rich user dialogues completed within the two months of deployment, 2,500 of which were expert-annotated for a consequent validation. The validation results show that CLARITY surpasses human-level performance in terms of the first-attempt routing precision, naturally requiring up to 3 times shorter duration of the consultation than with a human.
@@
Vladimir Shaposhnikov1,2  Aleksandr Nesterov1  Ilia Kopanichuk1,3 Ivan Bakulin1,3Egor Zhelvakov1Ruslan Abramov4 Ekaterina Tsapieva5Dmitry V. Dylov1Ivan Oseledets1 1AIRI  2Skoltech  3MIPT  4SberMedAI  5Sber Russia {shaposhnikov, nesterov, kopanichuk}@airi.net
@@ 1 Introduction
Medical dialogue systems are designed to facilitate clinical consultations, assist healthcare providers, and connect patients to a proper expertise. Unlike general-purpose conversational AI, these systems must be able to handle complex diagnostic reasoning, extended back-and-forth interactions, and domain-specific medical knowledge. Although recent advances in LLMs have significantly improved their conversational abilities, they are still fundamentally flawed for a real-world adoption in healthcare.
Despite all the efforts, modern patient routing approaches often involve unnecessary delays. Patients typically consult general practitioners (GPs) as an intermediary step before being referred to specialized experts, which increases wait times and complicates timely access to care. Automated assessment of patient conditions with a generated diagnostic hypotheses could significantly streamline this process, benefiting both patients and healthcare providers Nazi and Peng (2024). This optimization also promises substantial economic advantages, potentially reducing healthcare costs and improving the allocation of resources.
@@
Several specialized LLMs, including AMIE, DISC-MedLLM, CDD, and CoD Tu et al. (2024); Bao et al. (2023); Chen et al. (2024), demonstrate the potential of this technology. While systems like AMIE show high diagnostic accuracy, challenges remain regarding scalability and reliance on specialized datasets. Others, such as DISC-MedLLM, prioritize empathetic patient interaction, while CDD focuses on structured symptom questioning. However, the field lacks unified standards for training, evaluation, and deployment, hindering model comparison and regulatory approval Bedi et al. (2024); Singhal et al. (2023).
@@
To address these challenges, this article introduces CLARITY, a hybrid system integrating the strengths of rule-based and LLM approaches. CLARITY is designed to provide a flexible, controllable, and accurate medical dialog system suitable for real-world applications. By leveraging FSM, a microservices architecture, and specialized datasets, CLARITY facilitates structured, context-aware patient interactions, enhancing diagnostic reliability, enabling real-time critical condition recognition, and providing personalized patient guidance.
@@ 2 System Architecture
The CLARITY system is designed as a hybrid architecture, integrating a FSM for the management of dialogue with a microservices for the processing of requests.
@@
  • FSM governs the dialogue flow by managing states and transitions based on user input and context.
  • Text generation services are responsible for a natural language generation and complex query processing.
@@ 2.1 Chat Manager
States. The FSM has six dialogue contexts: Initialization—greets the user and begins collecting initial complaints; Information Collection—coordinates the information-gathering module for symptoms and medical history; Diagnosis—processes collected data to generate preliminary diagnoses, select medical specialty, and formulate recommendations; Moderation—prevents unsafe dialogue scenarios; Emergency—executes predefined protocols for urgent cases; and Free Dialogue—handles non-standard requests via the answer-generation module.
@@ 2.2.5 Information Collector
1. The module begins by receiving the user’s message and dialogue history. 2. It then searches the historical dialogue vector database; if a similar case is found (cosine similarity >0.965>0.965), previously generated questions are reused. 3. If no match is identified, the system uses a domain-specific prompt to generate five new candidate questions with the LLM. 4. To avoid redundancy, each candidate is compared to prior dialogue questions, and those with cosine similarity above 0.86 are discarded. 5. Remaining candidates are passed through a relevance classifier trained on 2,500 labeled examples; the classifier achieves 0.84 precision on the "relevant" class. 6. The most relevant question is selected and presented to the user. This architecture ensures both contextual accuracy and low response latency.
@@ 2.2.6 Medical Specialty Selector
1. Diagnosis Generation: Using the patient’s complaints and medical history, the system queries the LLM to generate $N$ possible diagnoses. If fewer than $N$ are returned, the request is repeated to ensure sufficient coverage.
2. Specialist Selection: For each diagnosis, a parallel LLM call determines the most relevant medical specialist (e.g., cardiologist for cardiovascular symptoms, gastroenterologist for digestive issues), ensuring that each condition is matched with an appropriate expert.
3. Explanation Generation: The system produces a short explanation for each diagnosis-specialist pair, including a description of the condition, reasoning behind the diagnosis, and rationale for the referral. These explanations aim to improve patient understanding and transparency.
@@ 2.2.7 Answer Generator
2. Once the standard dialogue script has been completed and all necessary information has been collected, as well as a preliminary diagnosis provided, the dialogue can be considered complete. In this instance, the open-dialogue module permits the user to pose supplementary queries, seek elucidation, or request general health-related information.
@@ 3 Experiments And Results
A comprehensive breakdown of datasets, model architectures and training protocols is available in Appendix D, while Tables 1 and 2 collate the key service-level metrics and pilot statistics across all four scenarios. Figure 2 reproduces a complete Transparency consultation, full transcripts for the remaining cases—Critical, Safety, and Adaptability—are available in Appendix C (Table 4).
@@ 3.1 First Pilot Study: Understanding User Behavior And Preferences
The pilot study resulted in 785 distinct preliminary diagnoses, with referrals made to 70 physicians across various specialties. The most frequently recommended specialists were neurologists, otolaryngologists, and gastroenterologists, which is indicative of the relevance of these specialties within the scope of the pilot study (Appx. K). Neurological, renal, and gastrointestinal disorders were the most commonly diagnosed conditions. A comparison of the socio-demographic characteristics of the participants and those of the overall user base of the platform indicated a correspondence between the two groups in terms of both age and gender (Appx. H and Appx. I).
A total of 18.2% of users indicated that they found the consultation to be fully useful, while 54.5% of respondents rated it as partially useful. However, 31% of users reported that the number of questions asked was insufficient, potentially pointing to limitations in audience targeting or the depth of the dialogue process. On a positive note, 80% of users rated the interaction as friendly, which is a key factor in maintaining user engagement and satisfaction (Appx. J).
@@ 3.2 Second Pilot Study: Focus On Real-World Application
A second pilot study was conducted to evaluate the real-world application of CLARITY in a larger-scale setting. This study focused less on audience behavior and more on the practical implementation and outcomes of the system. During the pilot period, a total of 55,856 dialogues were conducted. The dialogue funnel analysis revealed a 92.3% conversion rate to the next step after dialogue initiation, with 54.7% of dialogues resulting in a diagnosis and referral to a physician. In terms of appointment conversions, online consultations had a conversion rate of approximately 26.8%, while offline consultations had a conversion rate of approximately 5.6%.
@@ 4 Discussion
The system achieved remarkable performance in specialist routing, with 77% precision for first recommendations and 96% recall for top-3 recommendations. This surpasses typical human-level performance in initial routing while requiring only one-third of the usual consultation time. These results are particularly significant given current healthcare bottlenecks. The traditional model, where patients must first consult general practitioners for specialist referrals, creates substantial delays in accessing specialized care. CLARITY demonstrates that AI systems can effectively assume this initial triage function, potentially accelerating patient access to appropriate specialists while reducing the burden on primary care providers.
@@
The 80% positive interaction ratings demonstrate the system’s ability to maintain empathetic and professional communication while adhering to clinical protocols. Operational Efficiency: The mean consultation time of 2 minutes 13 seconds represents a significant improvement over traditional consultations, without compromising diagnostic accuracy.
@@ 6 Limitations
  • Undisclosed LLM identity. The paper does not specify which large language model was used due to anonymity constraints at this stage. While this limits reproducibility and benchmarking, the system is explicitly designed to be LLM-agnostic and compatible with any modern foundation model. Full technical details will be made available upon publication.
@@
  • Lack of clinical outcome validation. Although diagnostic and referral suggestions were quantitatively evaluated, long-term patient outcomes were not tracked. This is due to regulatory and data privacy limitations in pilot deployments. As part of ongoing collaborations with clinical partners, we plan to conduct retrospective chart reviews and, where possible, prospective outcome-linked evaluations in future studies.
@@
  • Partial reproducibility. Until the end of the anonymity period several code components and the production LLM name remain under NDA, so full replication is not yet possible. Upon acceptance we will release all prompt templates, training/evaluation scripts, and any source code that is not contract-restricted.
  • Lack of unified quantitative benchmarks. There is still no publicly available dataset or protocol that simultaneously covers triage, speciality selection, emergency detection, and safety filtering. Numerical head-to-head evaluation with earlier systems is therefore impossible without re-implementing each approach on private clinical data. As a partial remedy we performed a qualitative, functional comparison of CLARITY with seven representative systems (AMIE, DISC-MedLLM, CoD, Polaris, MedAgents, Dual-Inf, and other). The resulting matrix—Appendix A Table 3 — contrasts core capabilities and highlights CLARITY’s unique FSM-based architecture. While this analysis cannot replace a common benchmark, it offers readers a transparent, like-for-like view of where our system stands relative to prior art and exposes gaps that future community datasets should address.
@@ A.1 Evolution of Medical Dialogue Systems
    Task-Specific AI Models – Machine learning-based systems such as Babylon Health Moramarco et al. (2021) and Ada Health Kühnel et al. (2023) introduced data-driven triage and symptom assessment, but they remained single-turn and failed to engage in dynamic diagnostic transparency.
@@
    LLM-Powered Conversational Agents – Recent systems, such as AMIE Tu et al. (2024), DISC-MedLLM Bao et al. (2023), CoD Chen et al. (2024) and MEDAGENTS Tang et al. (2024), integrate structured multi-turn dialogue processing and medical reasoning, improving diagnostic accuracy and interaction depth.

### v2

@@ Abstract
We present CLARITY (Clinical Assistant for Routing, Inference and Triage), an AI-driven platform designed to facilitate patient-to-specialist routing, clinical consultations, and severity assessment of patient conditions. Its hybrid architecture combines a Finite State Machine (FSM) for structured dialogue flows with collaborative agents that employ Large Language Model (LLM) to analyze symptoms and prioritize referrals to appropriate specialists. Built on a modular microservices framework, CLARITY ensures safe, efficient, and robust performance, flexible and readily scalable to meet the demands of existing workflows and IT solutions in healthcare.
We report integration of our clinical assistant into a large-scale national interhospital platform, with more than 55,000 content-rich user dialogues completed within the two months of deployment, 2,500 of which were expert-annotated for subsequent validation. The validation results show that CLARITY surpasses human-level performance in terms of the first-attempt routing precision, naturally requiring up to 3 times shorter duration of the consultation than with a human.
@@
Vladimir Shaposhnikov*,1,2  Aleksandr Nesterov*,1  Ilia Kopanichuk*,1,3 Ivan Bakulin1,3Egor Zhelvakov1Ruslan Abramov4 Ekaterina Tsapieva5Iaroslav Bespalov1Dmitry V. Dylov1Ivan Oseledets1 1AIRI  2Skoltech  3MIPT  4SberMedAI  5Sber * Equal contribution Russia {shaposhnikov, nesterov, kopanichuk}@airi.net
@@ 1 Introduction
Medical dialogue systems are designed to facilitate clinical consultations, assist healthcare providers, and connect patients with proper expertise. Unlike general-purpose conversational AI, these systems must be able to handle complex diagnostic hypotheses reasoning, extended back-and-forth interactions, and domain-specific medical knowledge. Although recent advances in LLMs have significantly improved their conversational abilities, they are still fundamentally flawed for real-world adoption in healthcare.
Despite all the efforts, modern patient routing approaches often involve unnecessary delays. Patients typically consult general practitioners (GPs) as an intermediary step before being referred to specialized experts, which increases wait times and complicates timely access to care. Automated assessment of patient conditions with a generated diagnostic hypotheses could significantly streamline this process, benefiting both patients and healthcare providers Umerenkov et al. (2025); Nazi and Peng (2024). This optimization also promises substantial economic advantages, potentially reducing healthcare costs and improving the allocation of resources.
@@
Several specialized LLMs, including AMIE, DISC-MedLLM, CDD, and CoD Tu et al. (2024); Bao et al. (2023); Chen et al. (2024), demonstrate the potential of this technology. While systems like AMIE show high diagnostic hypotheses accuracy, challenges remain regarding scalability and reliance on specialized datasets. Others, such as DISC-MedLLM, prioritize empathetic patient interaction, while CDD focuses on structured symptom questioning. However, the field lacks unified standards for training, evaluation, and deployment, hindering model comparison and regulatory approval Bedi et al. (2024); Singhal et al. (2023).
@@
To address these challenges, this article introduces CLARITY, a hybrid system integrating the strengths of rule-based and LLM approaches. CLARITY is designed to provide a flexible, controllable, and accurate medical dialog system suitable for real-world applications. By leveraging FSM, a microservices architecture, and specialized datasets, CLARITY facilitates structured, context-aware patient interactions, enhancing diagnostic hypotheses reliability, enabling real-time critical condition recognition, and providing personalized patient guidance.
@@ 2 System Architecture
The CLARITY system is designed as a hybrid architecture, integrating an FSM for the management of dialogue with a microservices architecture for the processing of requests.
@@
  • FSM governs the dialogue flow by managing states and transitions based on user input and context. Text generation services are responsible for a natural language generation and complex query processing. In our deployment, the LLM used for generation and analysis is GigaChat [^1] https://giga.chat/.
@@ 2.1 Chat Manager
States. The FSM has six dialogue contexts:
  • Initialization - greets the user and begins collecting initial complaints;
  • Information Collection - coordinates the information - gathering module for symptoms and medical history;
  • Diagnostic Hypothesis & Routing - processes collected data to generate preliminary diagnostic hypotheses, select medical specialty, and formulate recommendations;
  • Moderation - prevents unsafe dialogue scenarios;
  • Emergency - executes predefined protocols for urgent cases;
  • Free Dialogue - handles non-standard requests via the answer-generation module.
@@
2.2 FSM Scalability
CLARITY’s FSM addresses scalability concerns through hierarchical decomposition designed to prevent architectural brittleness as clinical scope expands:
Multi-level FSM Design. The current implementation employs logical decomposition where a global FSM graph describes overall system behavior and manages high-level transitions between initialization, information collection, diagnostic hypotheses, moderation, and emergency handling states. Simultaneously, local FSM graphs govern the logic within individual medical scenarios and specialty-specific workflows, creating a hierarchical structure that naturally distributes complexity across multiple manageable components.
Modular Expansion Strategy. New medical domains can be integrated by:
  • Defining domain-specific local FSM graphs that handle specialty workflows.
  • Adding appropriate transition points in the global FSM to route users to new domains.
  • Implementing corresponding microservices that provide domain-specific processing capabilities.
This approach ensures system complexity grows linearly rather than exponentially with clinical area additions.
Graph-based Storage and Analysis. The FSM is stored as a graph structure in memory, enabling development of semi-automated analysis tools for: redundant cycle detection, unreachable state identification, conflicting transition condition analysis, and other architectural consistency checks that support long-term maintainability.
@@ 2.3.5 Information Collector
  1. 1.
    The module begins by receiving the user’s message and dialogue history.
  2. 2.
    It then searches the historical dialogue vector database; if a similar case is found (cosine similarity >0.965>0.965), previously generated questions are reused.
  3. 3.
    If no match is identified, the system uses a domain-specific prompt to generate five new candidate questions with the LLM.
  4. 4.
    To avoid redundancy, each candidate is compared to prior dialogue questions, and those with cosine similarity above 0.86 are discarded.
  5. 5.
    Remaining candidates are passed through a relevance classifier trained on 2,500 labeled examples; the classifier achieves 0.84 precision on the "relevant" class.
  6. 6.
    The most relevant question is selected and presented to the user. This architecture ensures both contextual accuracy and low response latency.
@@ 2.3.6 Medical Specialty Selector
1. Diagnostic Hypotheses Generation: Using the patient’s complaints and medical history, the system queries the LLM to generate $N$ possible diagnostic hypotheses. If fewer than $N$ are returned, the request is repeated to ensure sufficient coverage.
2. Specialist Selection: For each diagnostic hypotheses, a parallel LLM call determines the most relevant medical specialist (e.g., cardiologist for cardiovascular symptoms, gastroenterologist for digestive issues), ensuring that each condition is matched with an appropriate expert.
3. Explanation Generation: The system produces a short explanation for each diagnosis-specialist pair, including a description of the condition, reasoning behind the diagnostic hypotheses, and rationale for the referral. These explanations aim to improve patient understanding and transparency.
@@ 2.3.7 Answer Generator
2. Once the standard dialogue script has been completed and all necessary information has been collected, as well as a preliminary diagnostic hypotheses provided, the dialogue can be considered complete. In this instance, the open-dialogue module permits the user to pose supplementary queries, seek elucidation, or request general health-related information.
@@ 3 Experiments And Results
A comprehensive breakdown of datasets, model architectures and training protocols is available in Appendix D, while Tables 1 and 2 collate the key service-level metrics and pilot statistics across all four scenarios. Figure 2 reproduces a complete Transparency consultation, full transcripts for the remaining cases - Critical, Safety, and Adaptability - are available in Appendix C (Table 4).
@@ 3.1 First Pilot Study: Understanding User Behavior And Preferences
The pilot studies were conducted on a nationwide inter-hospital telehealth platform where CLARITY was fully integrated into the standard user onboarding and consultation workflow; users interacted with the system in their usual telehealth environment (both new registrants and returning users). Participants received clear onboarding notifications about the AI assistant and could request a hand-off to a human specialist at any time.
@@
The pilot study resulted in 785 distinct preliminary diagnostic hypotheses, with referrals made to 70 physicians across various specialties. The most frequently recommended specialists were neurologists, otolaryngologists, and gastroenterologists, which is indicative of the relevance of these specialties within the scope of the pilot study (Appendix K). Neurological, renal, and gastrointestinal disorders were the most commonly diagnosed conditions. A comparison of the socio-demographic characteristics of the participants and those of the overall user base of the platform indicated a correspondence between the two groups in terms of both age and gender (Appendix H and Appendix I).
A total of 18.2% of users indicated that they found the consultation to be fully useful, while 54.5% of respondents rated it as partially useful. However, 31% of users reported that the number of questions asked was insufficient, potentially pointing to limitations in audience targeting or the depth of the dialogue process. On a positive note, 80% of users rated the interaction as friendly, which is a key factor in maintaining user engagement and satisfaction (Appendix J).
@@ 3.2 Second Pilot Study: Focus On Real-World Application
A second pilot study was conducted to evaluate the real-world application of CLARITY in a larger-scale setting. This study focused less on audience behavior and more on the practical implementation and outcomes of the system. During the pilot period, a total of 55,856 dialogues were conducted. The dialogue funnel analysis revealed a 92.3% conversion rate to the next step after dialogue initiation, with 54.7% of dialogues resulting in diagnostic hypotheses and a referral to a physician. In all cases, CLARITY performed the initial triage and routed users toward the appropriate channel. For appointment conversions, “online” consultations (telemedicine appointments booked and conducted via the platform) converted at  26.8%, whereas “offline” consultations (in-person visits at clinics or hospitals) converted at  5.6%.
@@ 4 Discussion
The system achieved remarkable performance in specialist routing, with 77% precision for first recommendations and 96% recall for top-3 recommendations. This surpasses typical human-level performance in initial routing while requiring only one-third of the usual consultation time (mean 2m 13s vs $\approx$ 6-7 minutes for GP practice). These results are particularly significant given current healthcare bottlenecks. The traditional model, where patients must first consult general practitioners for specialist referrals, creates substantial delays in accessing specialized care. CLARITY demonstrates that AI systems can effectively assume this initial triage function, potentially accelerating patient access to appropriate specialists while reducing the burden on primary care providers.
@@
The 80% positive interaction ratings demonstrate the system’s ability to maintain empathetic and professional communication while adhering to clinical protocols. Operational Efficiency: The mean consultation time of 2 minutes 13 seconds represents a significant improvement over traditional consultations, without compromising diagnostic hypotheses accuracy.
@@ 5 Conclusion
6 Contribution Statement
Vladimir Shaposhnikov, Aleksandr Nesterov, and Ilia Kopanichuk contributed equally: they formulated the research, designed and ran the experiments, performed the analysis, and drafted the core manuscript text. Ekaterina Tsapieva and Ruslan Abramov led the productization effort and guided the introduction of the solution into a real-world workflow. Ivan Bakulin and Egor Zhelvakov were responsible for technical deployment and production integration. Iaroslav Bespalov, Dmitry V. Dylov and Ivan Oseledets are co-senior authors who supervised the project and analysed the results. All authors took part in preparing the final manuscript.
@@ 7 Limitations
  • Model identity disclosure. The system uses GigaChat as the foundation model for generation and analysis. While this work is designed to be LLM-agnostic, all reported results were obtained with GigaChat under the production configuration used in deployment.
@@
  • Time–Satisfaction Mismatch and Explanation Factuality Limit We acknowledge that shorter consultations do not necessarily translate into higher user satisfaction or trust: although median time decreased, many users still rated interactions as only partially useful, and we did not establish a causal link between duration, perceived usefulness, and trust. In addition, while we evaluated routing accuracy, we did not conduct a dedicated factuality (hallucination) audit of generated explanations. Future iterations will include targeted user-trust studies and clinician-in-the-loop audits of explanation correctness.
  • Lack of clinical outcome validation. Although diagnostic hypotheses and referral suggestions were quantitatively evaluated, long-term patient outcomes were not tracked. This is due to regulatory and data privacy limitations in pilot deployments. As part of ongoing collaborations with clinical partners, we plan to conduct retrospective chart reviews and, where possible, prospective outcome-linked evaluations in future studies.
  • Absence of active learning pipeline. Currently, CLARITY does not implement real-time self-updating mechanism or active learning. This design choice reflects critical regulatory and ethical constraints guiding the deployment of medical AI. Medical AI systems must pass stringent validation before algorithmic changes can be implemented, making real-time updates impossible from a regulatory standpoint. Furthermore, self-updating models based on patient interactions raise privacy concerns and may lead to unpredictable behavioural changes that undermine confidence in clinical decision-making. As future work, we are developing a controlled active learning pipeline that will operate under strict supervision and include offline batch learning cycles with mandatory clinical validation, explicit patient consent for the use of anonymised data, and differential privacy techniques.
@@
  • Partial reproducibility. We are committed to open-sourcing all available components of our system. The code will be publicly released at: https://github.com/AnonymousSubmission996/CLARITY upon acceptance. Certain production-specific services remain under NDA and cannot be distributed.
  • Lack of unified quantitative benchmarks. There is still no publicly available dataset or protocol that simultaneously covers triage, speciality selection, emergency detection, and safety filtering. Numerical head-to-head evaluation with earlier systems is therefore impossible without re-implementing each approach on private clinical data. As a partial remedy we performed a qualitative, functional comparison of CLARITY with seven representative systems (AMIE, DISC-MedLLM, CoD, Polaris, MedAgents, Dual-Inf, and other). The resulting matrix - Appendix A Table 3 - contrasts core capabilities and highlights CLARITY’s unique FSM-based architecture. While this analysis cannot replace a common benchmark, it offers readers a transparent, like-for-like view of where our system stands relative to prior art and exposes gaps that future community datasets should address.
@@ 8 Ethics
9 Acknowledgments
We are grateful to SberHealth for supporting the project and to all participating physicians for their invaluable feedback during development and evaluation. We also thank our clinical partners and annotators for their assistance with data labeling, which was essential for validating and improving the study’s outcomes.
@@ A.1 Evolution of Medical Dialogue Systems
    Task-Specific AI Models – Machine learning-based systems such as Babylon Health Moramarco et al. (2021) and Ada Health Kühnel et al. (2023) introduced data-driven triage and symptom assessment, but they remained single-turn and failed to engage in dynamic diagnostic hypotheses transparency.
@@
    LLM-Powered Conversational Agents – Recent systems, such as AMIE Tu et al. (2024), DISC-MedLLM Bao et al. (2023), CoD Chen et al. (2024) and MEDAGENTS Tang et al. (2024), integrate structured multi-turn dialogue processing and medical reasoning, improving diagnostic hypotheses accuracy and interaction depth.